OpenStudy (anonymous):

How many liters of peroxide should be added to 12 liters of an 8% peroxide solution so that the remaining solution contains 16% peroxide?

OpenStudy (anonymous):

lets say you add \(x\) liters of peroxide then the the total number of liters will be \(12+x\) presently you have 8% of 12 or \(.96\) liters of peroxide when you add \(x\) liters you will have \(.96+x\) liters of peroxide total you want that to be 16% of (12+x\) so set \[.96+x=.16(12+x)\] and solve for\(x\)

OpenStudy (anonymous):

probably easier to solve if you get rid of decimal by multiplying both sides by 100 and write \[96+100x=16(12+x)\]

OpenStudy (anonymous):

thank you